OpenStudy (anonymous):

A card is selected from a deck of 52 playing cards. Find the probability of selecting · an even number ten or less given the card is a not a diamond. · an Ace, given that the card is black. · a spade given the card is red. Show step by step work. Give all solutions exactly in reduced fraction form.

OpenStudy (anonymous):

Well, the first step is to find your total number of desired outcomes. Are you familiar with a deck of cards?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well simple probability is desired outcomes over total number of outcomes. How many even numbers are there, including 10?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

Ok, but we're taking out one of the faces...so no diamonds. How many desired outcomes do we have now?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

Ok, so your probability is going to be a fraction 15/52 because it is the desired outcome (15 cards are even and NOT a diamond) and 52 is the total number of cards. Can you simplify 15/52?

OpenStudy (anonymous):

what about the rest of the question. Can help with the other parts

OpenStudy (anonymous):

ok for the second one, how many total aces are there in a deck?

OpenStudy (zarkon):

why divide by 52?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

@Zarkon - because you express probability as [desired outcomes]/[total outcomes]. There are 52 cards total in a a deck.

OpenStudy (zarkon):

it is not 52 in this case

OpenStudy (anonymous):

@black12 ok so there are 4 aces. How many are going to be black aces?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

You got it! so now put in in fraction form: There are 2 black aces and 52 total cards. What is the fraction?

OpenStudy (anonymous):

2/52=1/26

OpenStudy (zarkon):

that is not correct either

OpenStudy (zarkon):

these are conditional probabilities...

OpenStudy (anonymous):

Zarkon, this is simple probability.

OpenStudy (zarkon):

doesn't seem that way...since I have yet to see a correct answer

OpenStudy (anonymous):

@black12 What grade are you in, if you don't mind me asking?

OpenStudy (anonymous):

Zarkon, I didn't mean that this is simple in a sarcastic way. I'm just saying that there is only one even going on. Would you care to explain your way of doing it?

OpenStudy (anonymous):

even = event

OpenStudy (anonymous):

COLLEGE

OpenStudy (zarkon):

again this is conditional probability.. @MissMai I am not being rude

OpenStudy (zarkon):

I'm giving you a fact

OpenStudy (anonymous):

zarkon can u show me ur way please

OpenStudy (zarkon):

\[P(A|B)=\frac{P(A\text{ and }B)}{P(B)}\]

OpenStudy (zarkon):

A=an even number ten or less B=card is a not a diamond.

OpenStudy (anonymous):

can you input it cause Iam leaning that but its confusing

OpenStudy (zarkon):

P(A and B)=15/52 P(B)=39/52

OpenStudy (zarkon):

P(A and B)/P(B)=15/39

OpenStudy (anonymous):

I honestly think that it is simple probability because you are only selecting one card.

OpenStudy (zarkon):

"an even number ten or less \(\underline{\text{given}}\) the card is a not a diamond." it is conditional probability

OpenStudy (anonymous):

Explain how, please.

OpenStudy (zarkon):

the word 'GIVEN' tells us that this is a conditional probability problem

OpenStudy (anonymous):

b12 are you getting this or do you need it broken down?

OpenStudy (anonymous):

@Z - I get it now. It's been a while.